OpenStudy (anonymous):

How to integrate e^(e^x). I used taylor series to find the approximation. Is there another way of solving it? Thx

OpenStudy (anonymous):

Thanks mate but what do they mean by E?

OpenStudy (anonymous):

And how did it arrive to that answer

OpenStudy (kainui):

The only way I can see to integrate it is this way: \[e^y = \sum_{n=0}^\infty \frac{y^n}{n!}\] \[e^{e^x} = \sum_{n=0}^\infty \frac{e^{xn}}{n!}\] now you can integrate term by term of this power series.

OpenStudy (anonymous):

Could you please show me the whole solution?

OpenStudy (kainui):

Nah, there's not really unless you want to define a new function.